find the center and the radius of the circle x2+y2-10x-6y+29=0. Draw the graph
Answer by Nate(3500)

x^2 - 10x + y^2 - 6y = -29
(x - 5)^2 + (y - 3)^2 = -29 + 25 + 9 = 5
c(5,3) and r = sqrt(5)